# Axiomatic Causal Interventions for Reverse Engineering Relevance Computation in Neural Retrieval Models

### v1

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657841isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and rankingccs: Information systems Document representation
@@ 5.1. Importance of Added/Deleted Query Terms
First, we find that the model becomes confident and reaches a decision in the later layers, specifically in Layers 4 and 5 (Figure 4). At this point, term frequency information transfers from the existing selected query term tokens $tok_{qterm^{+}}$ to the CLS token. This shift towards the CLS token is expected, given that the ranking score is derived from a pooled representation of the CLS token.
Second, for TFC1-I, the injected tokens ($tok_{inj}$) surprisingly are not the most important for recovering performance. Rather, the instances of the selected query term already present in the original document ($tok_{qterm^{+}}$) are the most impactful. We postulate that the model may store important information in query terms situated toward the beginning of the document. To further investigate this hypothesis, we run an additional experiment changing the location of the perturbation, injecting the selected query term at the beginning of the document rather than the end. By doing so, we find that this leads to a shift in importance towards the injected tokens at the beginning of the document (Figure 5), suggesting that the model stores the majority of the term frequency information in the initial occurrence of duplicate terms.
@@ 6.2. Relevance Computation Heads
Activation patching on individual attention heads reveals four heads that significantly express a term-frequency signal aligned with the TFC1 axiom. However, as previously noted in Section 5, inspecting the attention scores and patching along token positions reveals distinct behavior among the heads. Specifically, heads in earlier layers (i.e., 0.9 and 1.6) function as duplicate token heads, primarily attending to duplicate instances of the selected query term (thereby potentially counting term frequencies) and storing important relevance information in these tokens. Remarkably, these heads can recover a significant amount of ranking performance on duplicate tokens alone, sometimes even exceeding the original performance, indicating that the model encapsulates a robust relevance signal in these tokens (Figure 10).
On the other hand, heads in middle layers (i.e., 2.3 and 3.8) exhibit distinct behavior compared to heads in earlier layers. These middle-layer heads do not attend to duplicate tokens, yet they still have a strong positive impact on performance. Furthermore, while patching the entire head significantly influences model performance in a positive direction, no one single token type is responsible for the majority of this behavior, suggesting that the term frequency signal may not be concentrated in any individual component (Figure 10).
@@
Figure 10. Activation patching results by token position for attention heads 0.9, 1.6, 2.3, and 3.8 for documents with duplicate occurrences of the selected query term. Heads in earlier layers increase in score by more than 75% just by patching on duplicate instances of the selected query term. On the other hand, heads in middle layers recover less than half (or even none at all) of the ranking score on a single token type.

### v2

@@ Abstract.
ccs: Information systems Retrieval models and rankingccs: Information systems Document representation
@@ 5.1. Importance of Added/Deleted Query Terms
First, we find that the model becomes confident and reaches a decision in the later layers, specifically in Layers 4 and 5 (Figure 4). At this point, term frequency information transfers from the injected tokens ($tok_{inj}$) and the existing selected query term tokens $tok_{qterm^{+}}$ to the CLS token. This shift towards the CLS token is expected, given that the ranking score is derived from a pooled representation of the CLS token.
Second, for TFC1-I, the injected tokens ($tok_{inj}$) surprisingly are not the only important for recovering performance. Rather, the instances of the selected query term already present in the original document ($tok_{qterm^{+}}$) are also impactful. We postulate that the model may store important information in query terms situated toward the beginning of the document. To further investigate this hypothesis, we run an additional experiment changing the location of the perturbation, injecting the selected query term at the beginning of the document rather than the end. By doing so, we find that this leads to a full shift in importance towards the injected tokens at the beginning of the document (Figure 5), suggesting that the model stores the majority of the term frequency information in the initial occurrence of duplicate terms.
@@ 6.2. Relevance Computation Heads
Activation patching on individual attention heads reveals four heads that significantly express a term-frequency signal aligned with the TFC1 axiom. However, as previously noted in Section 5, inspecting the attention scores and patching along token positions reveals distinct behavior among the heads. Specifically, heads in earlier layers (i.e., 0.9 and 1.6) function as duplicate token heads, primarily attending to duplicate instances of the selected query term (thereby potentially counting term frequencies) and storing important relevance information in these tokens (Figure 8). Interestingly, these heads can recover a significant amount of ranking performance on duplicate tokens alone, indicating that the model encapsulates a robust relevance signal in these tokens (Figure 10).
On the other hand, heads in middle layers (i.e., 2.3 and 3.8) exhibit distinct behavior compared to heads in earlier layers (Figure 8). These middle-layer heads do not attend to duplicate tokens, yet still have a strong positive impact on performance when patching across all tokens (Figure 6). Furthermore, while patching the entire head significantly influences model performance, no single token type is responsible for the majority of this behavior, suggesting that the term frequency signal may not be concentrated in any individual component (Figure 10).
@@
Figure 10. Activation patching results by token position for attention heads 0.9, 1.6, 2.3, and 3.8 for documents with duplicate occurrences of the selected query term. Heads in earlier layers increase in score just by patching on a single token type ($tok_{inj}$ for 0.9 and $tok_{qterm+}$ for 1.6). On the other hand, heads in middle layers exhibit varied behavior, with a single token type causing a decrease in score ($tok_{qterm+}$ for 2.3) and multiple token types causing an increase in score ($tok_{CLS}$ and $tok_{qterm+}$ for 3.8. These differences in behavior, combined with the attention patterns, suggest that the middle-layer heads may be more sensitive to information from across the entire document rather than focusing on individual tokens.
@@ Acknowledgements.
Appendix A Author’s Note
A version of this paper appeared in SIGIR 2024. Since publication, we have found a bug in our plotting code that affected the visualization of some patching results in the original paper. The bug has since been corrected, and the code to reproduce all experiments and plots has been updated in the codebase [^2] https://github.com/catherineschen/axiomatic-ir-interventions. This version of the paper includes all corrected figures and updated captions and references, and for complete transparency, we list and provide more context to all updates below.
We note that the main conclusion of the paper, that activation patching can be used to isolate behavior to specific components and tokens in neural retrieval models, remains unchanged. We apologize for any confusion this may have caused and appreciate your understanding as we work to maintain transparency and reproducibility in all of our research.
A.1. Updates to Figures and Text
Figures 4 and 5
In Section 5.1, the second sentence in the second paragraph now reads (updates are bolded) “At this point, term frequency information transfers from the injected tokens ($tok_{inj}$) and the existing selected query term tokens $tok_{qterm^{+}}$ to the CLS token.”. This update reflects corrections to the left plot in Figures 4 and 5. Additionally, the last sentence in the third paragraph now reads “By doing so, we find that this leads to a full shift in importance towards the injected tokens at the beginning of the document (Figure 5)…”
Overall, both experiments in Figure 5 still indicate a positional bias towards the beginning of the document. When a query term is added to the end of a document, the injected term stores important information, but the initial occurrences of duplicate terms also hold significance. On the other hand, when a query term is added to the beginning, only the initial instance of the term retains importance. The consistent emphasis on initial occurrences across both perturbation locations is an interesting pattern that could be explored further in future research.
Figure 6
Figure 6 is largely consistent with the previous version, with the main difference being a slight reduction in the magnitude of some cell values. Nevertheless, the results continue to highlight that the same four attention heads (0.9, 1.6, 2.3, 3.8) have a significant effect when patched.
Figure 10
The caption for Figure 10 previously read “Activation patching results by token position for attention heads 0.9, 1.6, 2.3, and 3.8 for documents with duplicate occurrences of the selected query term. Heads in earlier layers increase in score by more than 75% just by patching on duplicate instances of the selected query term. On the other hand, heads in middle layers recover less than half (or even none at all) of the ranking score on a single token type.” and now reads “Activation patching results by token position for attention heads 0.9, 1.6, 2.3, and 3.8 for documents with duplicate occurrences of the selected query term. Heads in earlier layers increase in score just by patching on a single token type ($tok_{inj}$ for 0.9 and $tok_{qterm+}$ for 1.6). On the other hand, heads in middle layers exhibit varied behavior, with a single token type causing a decrease in score ($tok_{qterm+}$ for 2.3) and multiple token types causing an increase in score ($tok_{CLS}$ and $tok_{qterm+}$ for 3.8. These differences in behavior, combined with the attention patterns, suggest that the middle-layer heads may be more sensitive to information from across the entire document rather than focusing on individual tokens.”
While the updated Figure 10 reflects some changes, the distinction between heads in earlier layers vs heads in middle layers is less apparent from the patching results alone. However, the difference in the behavior of these heads can still be observed in their attention patterns (Figure 8): heads in earlier layers tend to focus on individual tokens, while heads in later layers have information dispersed across the entire document. This supports the primary aim of our paper, which is to demonstrate how activation patching can reveal finer-grained insights into the token interactions and attention head behavior that contribute to determining relevance.
As discussed in Section 6, we view this work as a first step toward gaining a deeper understanding of the internal mechanisms that neural retrieval models use to compute relevance. Our findings highlight the potential of activation patching for more granular analysis than what was previously possible. We hope future research can build upon our preliminary study and explore meaningful directions for further investigation into these mechanisms.